A camera is mounted 3.000 feet from a rocket launching pad. The camera needs to swivel as the ro launched to keep it in focus, a) If the rocket is rising vertically at 800 ft/sec when it is 4,000 feet hi fast is the camera-to-rocket distance changing?


1
Expert's answer
2021-02-01T07:25:35-0500

The camera-to-rocket distance as a function of time t is

"s(t)=\sqrt{3000^2+(800t)^2}=\sqrt{3000^2+800^2t^2}"


"s'(t)=\frac{1}{2}\frac{800^2*2t}{\sqrt{3000^2+800^2t^2}}="

"=\frac{800^2t}{\sqrt{3000^2+800^2t^2}}"


when rocket is 4,000 feet hi "t=\frac{4000}{800}=5"


"s'(5)=\frac{800^2*5}{\sqrt{3000^2+800^2*5^2}}="

"=\frac{640000*5}{\sqrt{9000000+16000000}}="

"=\frac{640000*5}{\sqrt{25000000}}=\frac{640000*5}{5000}=640"


Answer: the camera-to-rocket distance is rising with velocity 640 feet/sec.